OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4> 20.3° 10.2° 0.2° 30.3°

OpenStudy (cwrw238):

first find the scalar product of the 2 vectors can you do this?

OpenStudy (anonymous):

i do not remember well

OpenStudy (jdoe0001):

$$ cos(\theta)\cfrac{|u|\times |v|}{||u||\times ||v||} $$

OpenStudy (cwrw238):

its also called the dot product u.v = 6*7 + -1*-4 = 42 + 4 = 46 you find the angle (x) using the formula cos x = u.v ---- |u|.|v| where |u| and |v| are the magnitudes of the 2 vectors

OpenStudy (jdoe0001):

$$ \color{blue}{|u|\times |v|} \implies < a, b > \times < c , d >\\ \implies (a\times c) + (b \times d)\\ \color{blue}{||u|| \times ||v||} \implies \text{magnitudes of each multiplied}\\ \implies ||u|| = \sqrt{a^2+b^2}\\ \implies ||v|| = \sqrt{c^2+d^2} $$

OpenStudy (anonymous):

ok thanks i got it

OpenStudy (jdoe0001):

$$ cos(\theta)=\cfrac{|u|\times |v|}{||u||\times ||v||} $$ so you get cos() = SOMEVALUE arcsin both sides, to get the angle :)